OpenStudy (anonymous):

write quiotent as a rational expression and simplify 2y+10/3y-15 divided by 7y-21/8y-40

OpenStudy (anonymous):

Invert quotient and multiply instead. (2y+10)(8y-40) / (3y-15)(7y-21) =2(y+5)8(y-5) / 3(y-5)7(y-3) =16(y+5) / 21(y-3)